OpenStudy (anonymous):

Parametrize (x-3)^2 + y^2 = 1 Thank you! :)

OpenStudy (chrisasl):

In polar coordinates: x and y are equal to?

OpenStudy (anonymous):

x= rcost y=rsint

OpenStudy (anonymous):

Still can't find the parametrization

OpenStudy (chrisasl):

Sorry, can't think of the solution now :-/ If I think of something, I'll reply. Good luck with this!

OpenStudy (anonymous):

Thank you! :) I'll ask my professor tomorrow, I'll post the solution once I have it, but I'll try to keep figuring it out

OpenStudy (chrisasl):

We could do that:\[x^2 + y^2 - 6*x + 9 = 1 \] Or: \[x^2 + y^2 - 6*x = -8\] after that we plug in the formulas for x and y, so: \[r^2 - 6*r*\cos(t) = -8\] But I'm not sure that this is correct. Do as you said and reply with the solution after you ask your professor.

OpenStudy (anonymous):

ok so, I think I found it out :) \[r(t) = 3+\cos(t), \sin(t)\]

OpenStudy (chrisasl):

Hmm could you tell me the way?